OpenStudy (anonymous):

Which inequality will have a shaded area above its graph? 2x − y ≥ 4 4x + 3y < 6 x − 9y ≤ 1 x − 3y > 4

OpenStudy (michele_laino):

I think is x-9y <= 1, because if you solve that inequality for y, you get: \[y \ge \frac{ x-1 }{ 9 }\] which has the characteristic required

OpenStudy (michele_laino):